Find all solutions for (Let @ represent theta)
(2sin@-1)(2cos@+3)=0
Thank you in advance for your help!

Answer by Alan3354(69443)

Find all solutions for (Let @ represent theta)
(2sin@-1)(2cos@+3)=0
---------------
Either 2sin - 1 = 0
sin = 1/2
theta = 30, 150 + 2n*360, n = 0, 1, 2, 3... degs
---
or
2cos + 3 = 0
cos = -3/2 No real number solutions